# Automating Knowledge Discovery from Scientific Literature via LLMs: A Dual-Agent Approach with Progressive Ontology Prompting

### v1

@@ Introduction
In the LLM-Duo framework, the explorer is a chatbot built on LLM with RAG using Llamaindexhttps://www.llamaindex.ai. We use OpenAI ‘text-embedding-3-large’ as the embedding model and set the chunk size to 256 tokens with an overlapping size of 128. Particularly, we use ‘FastCoref’ (Otmazgin, Cattan, and Goldberg 2022) to process text chunks for coreference resolution before text embedding. Besides, we pass the document ID as metadata and apply a metadata filter on the chat engine to ensure the explorer only answers within the text of the specific document to be annotated. We set the retrieval to be on the top 8 text chunks based on similarity scores reranked with SentenceTransformerRerankhttps://docs.llamaindex.ai/en/stable/examples/node˙postprocessor/SentenceTransformerRerank employing the ‘cross-encoder/ms-marco-MiniLM-L-2-v2’https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-2-v2 model in Llamaindex. The evaluator is an external LLM who does not share any document context with the explorer.
Case Description: Speech-language Intervention Discovery
Speech-language therapy provides interventions for individuals with speech-language deficits, enhancing their quality of life across various life stages. When choosing an intervention, evidence-based practice (EBP) is attractive as it integrates research evidence from literature into the decision-making process to ensure high-quality patient care (Law et al. 1996). Intervention research, especially studies that offer clear intervention frameworks and comprehensive case studies, are valuable references to guide EBP designs. Intervention discovery aims to extensively gather speech-language interventions from the literature corpus as references to facilitate EBP design. It involves identifying relevant studies and extracting essential features of interventions including target disorder, procedure, efficacy, case study, therapy activity, etc., which is extremely labor-intensive for human reviewers, highlighting the efficiency of automating knowledge discovery based on LLMs.
To verify the effectiveness of our method in a realistic scenario, we employ our framework in a speech-language intervention discovery setting. The ontology is shown in Figure 4. We list the concept and relation definitions of this ontology in Appendix 1.1. To enable a large-scale discovery, we cultivate a literature base including 64,177 papers within the domain of speech-language therapy. Specifically, we employ a set of keywords carefully selected from a glossary of commonly used terms in speech-language therapy to guide our literature search. The search keywords and literature base are shown in Appendix 1.2.
@@
Baselines
The basic idea of LLM-Duo is to make LLM rectify its initial annotation results through external feedback from another LLM. Instead of setting up another LLM for evaluation, recent studies show that LLMs can perform self-correction to refine their responses by themselves (Liu et al. 2024; Li et al. 2024), and successful examples include Chain-of-Thought (CoT) (Wei et al. 2022) and SelfRefine (Madaan et al. 2024). We separately equip the chatbot based on RAG with these two prompting methods for annotation and denote them as CoT-RAG and SelfRefine-RAG. The full prompts for these two methods can be found in Appendix 2. Additionally, in LLM-Duo, a potential substitution of explorer is long-context LLM, which is capable of processing entire document tokens instead of chunking and retrieval with RAG. We denote LLM-Duo respectively using explorers built on RAG and long-context LLM as LLM-Duo-RAG and LLM-Duo-LongContext. Besides, we also compare with methods that directly input paper text to LLMs for zero-shot QA annotation without an evaluation feedback loop, including ShortContext LLM, LongContext LLM, OpenAI Assistant, and RAG.
(a) Illustration of participant annotation under different k values.
(b) Range of participant annotation accuracy using LLM-Duo-RAG with different LLMs at k=1,2,3.
(c) Token count distribution of retrieved-back chunks across varying similarity scores under participant annotation queries of k=1,2,3.
Figure 5: Overview of participant annotation and related metrics across different k values.
Evaluation
In our experiment of comparing with baseline approaches, we report six types of metrics: 1) Consistency Rounds (CR): the number of refine loops the method makes before achieving the annotation consistency; 2) Verbosity Index (VI): the number of aspects per 1k tokens of the annotations, which is an important metric for annotation emphasizing content completeness; 3) Enumeration Quantity (EQ): the number of items listed in the annotations (i.e., therapy activities, therapy goals in our test case.); 4) Faithfulness (Faith): the extent of the annotation faithful to the provided text, which is measured by FaithfulnessEvaluatorhttps://docs.llamaindex.ai/en/stable/examples/evaluation/faithfulness˙eval of Llamaindex. 5) Accuracy (ACC): the percent of correct annotations in all LLM-provided annotations. 6) Cover: the percent of correct LLM-provided annotations to the total mentioned concept entities in the provided text.
Results
In this section, we first provide a detailed evaluation of our progressive ontology prompting algorithm and the LLM-Duo annotation framework. Then, we showcase the results of speech-language intervention discoveries using our automated knowledge discovery framework, which further demonstrates the effectiveness of our framework in addressing practical scientific knowledge discovery tasks.
Context Size Analysis
In our progressive ontology prompting algorithm, when composing the annotation prompt for a concept node, the knowledge discoveries and the local ontology structure within the context (visited nodes within $k$-hop neighborhood) of a specific concept node serve as a prefix to provide conditions for annotation questions. The $k$ value determines the diversity and volume of information provided. To evaluate how the context size of the POP algorithm affects annotation quality, we conducted experiments by applying different $k$ values to generate prompts for LLM-Duo-RAG annotation. As shown in Figure 5(a), we selected an ontology substructure related to the ‘participant’ concept for the experiment, which was based on a random selection of 8 speech-language therapy papers.
@@ Results
Table 1: Comparison with baseline methods using various base LLMs.
@@
Figure 6: Ablation study and discovered speech-language intervention statistics.
The resulting annotation accuracy of the participant concept is shown in Figure 5(b). The results indicate that as the context hop $k$ increases, annotation accuracy improves significantly, suggesting that a larger context provides more informative prompts, thereby enhancing annotation quality. Moreover, GPT-4-turbo consistently outperforms GPT-3.5-turbo across all $k$ values, demonstrating that more advanced language models can further improve annotation accuracy. Besides, we inspect the text chunks retrieved back by different annotation queries based on various $k$ values. We report the range of similarity scores and token count distribution of retrieved-back chunks in Figure 5(c). The similarity score represents the semantically relevancy between retrieved texts to annotation queries. From the results, we can observe that when $k=1$, the retrieved text chunks generally exhibit low similarity with the query, and the token count tends to decrease as the similarity score increases, leading to poor annotation quality. In contrast, larger $k$ values, particularly $k=2$, result in more relevant content being retrieved. For $k=2$, the token count shows an increasing trend with higher similarity scores, indicating that more substantial and relevant content is being captured, which results in better annotation quality.
LLM-Duo with Baseline Annotation Methods
@@
To evaluate the performance of our LLM-Duo annotation framework, we compare it with baseline methods in the context of speech-language intervention discovery. We randomly pick 8 papers from our speech-language literature base for annotation using various baselines. The comparison focuses on three key dimensions: 1) Intervention Recognition (IR), identifying intervention entities within the paper; 2) Intervention Aspect Summary (IAS), annotating an aspect (e.g., procedure, therapy activity, therapy goals) of an intervention approach, which requires capture and summarize all related information fragments from the paper; and 3) Intervention Knowledge Completion (IKC), linking interventions to theme classes (e.g., speech awareness, speech articulation, comprehension, foundation skills, etc.) and setting concept nodes (e.g., home, healthcare facilities, early childhood centers, schools, teletherapy, etc.). For IR and IKC tasks, we ask human annotators to complete annotation as golden results for comparison. For the IAS task, due to the individual bias of human interpretation, we only ask human annotators to tag the text fragments in the paper pertaining to a specific intervention aspect.
@@
The experimental results are reported in Table 1. It should be noted that we implemented ‘ShortContext’ using Llama3-instruct-70b (FP16) and Mistral-instruct-8x22b models (INT8). However, directly prompting these models with full paper text fails to produce annotations in a zero-shot QA setting. Their generations do not align with the annotation questions. From the results in Table 1, we observe that LLM-Duo-RAG outperforms all baselines. Even though GPT4-turbo has a 128k context window length and is capable of generating annotations, its annotation coverage remains inadequate. Integrating it with the LLM-Duo framework can significantly improve both the accuracy and the comprehensiveness of the annotations. Additionally, compared with simple RAG, self-correct prompting methods such as CoT and SelfRefine can significantly enhance annotations, but their performance is still worse than LLM-Duo-RAG. Instead of utilizing costly GPT models, LLM-Duo-RAG, which employs open-source models including Llama3-instruct-70b and Mistral-instruct-8x22b, can achieve comparable annotation quality.
Ablation Study
@@
In our framework, the chatbot explorer uses the RAG technique. We enhance generation quality with ’FastCoref’ for coreference resolution and rerank retrieved chunks by similarity score using the ’cross-encoder/ms-marco-MiniLM-L-2-v2 model’. This section presents ablation studies for both components. Since recognizing the intervention entity is the critical first step following the POP algorithm, we report the accuracy of intervention recognition in this study. As shown in Figure 6(a), the results demonstrate that removing these components significantly decreases annotation accuracy, showing the necessity of each module.
Speech-Language Intervention Discovery
@@
Our knowledge discovery framework has identified 2,421 interventions supported by case studies from 64,177 research papers in the field of speech-language therapy. The statistics of discovered interventions are presented in Figure 6(b) and Figure 6(c). A complete intervention annotation example is available in Appendix 3. 19 clinicians and students reviewed our annotations through online Google forms. In checking the intervention annotation and reading the corresponding literature, they reported that 89% of our annotations are valid and complete. Based on the intervention ontology in Figure 4 and annotations, we have constructed an intervention knowledge graph that will be publicly accessible. This knowledge graph is expected to be a valuable resource for speech-language domain experts, supporting evidence-based clinical decision-making, question-answering and recommendation, ultimately enhancing healthcare outcomes.
Conclusion
@@
In this paper, we developed a novel automatic knowledge discovery framework based on LLMs, which is characterized by a progressive ontology prompting algorithm and a dual-agent annotation system. The proposed method achieves superior performance compared with advanced baselines, enabling more accurate knowledge discovery. Applied to speech-language intervention discovery, this framework also curates a valuable, accessible intervention knowledge base for the speech-language therapy community.
Appendix A Appendix
1 Intervention Ontology and Literature
1.1 Intervention Ontology The intervention ontology is shown in Figure 4 of the main paper. In this section, we provide a detailed explanation of the concepts introduced in the ontology as follows:
  • Disorder represent the type of disorder that causes difficulties in an individual’s voice, speech, language, or swallowing functions.
  • Efficacy represents the conclusion about the effectiveness of intervention.
  • Case Study represents a detailed examination of the intervention on a particular individual or group with communication disorders. The purpose of a case study is to provide a deep understanding of the patient’s unique needs and assess the intervention effectiveness.
  • Age represent the age of experiment participant or claimed target population of the intervention. The age is quantified with a granularity of half a year. We additionally convert age to age groups including “newborn, infants, toddlers, children, preschoolers, school-age children, older children, youth, teens, adolescents, adult, young adult, middle aged, aged, senior”. Each specific age may be associated with multiple age groups. For instance, an individual aged 13 years could be categorized to the ‘teens,’ ‘adolescents,’ and ‘children’ age groups.
  • Language represent the speaking language of experiment participant in the case study of the intervention.

### v2

@@ 4.1 Implementation
For LLM-Duo, the explorer is a chatbot built on LLM with RAG, implemented with Llamaindex [^2] https://www.llamaindex.ai framework. We use OpenAI ‘text-embedding-3-large’333 https://platform.openai.com/docs/guides/embeddings/embedding-models as the embedding model and set the chunk size to 256 tokens with an overlapping size of 128. Particularly, we use ‘FastCoref’ Otmazgin et al. (2022) to process text chunks for coreference resolution before text embedding. Additionally, we include the document ID as metadata for chunks and apply a metadata filter in the chat engine to ensure that the explorer only answers based on the specific document being annotated. We use Chroma [^4] https://github.com/chroma-core/chroma as the vector database. We set the retrieval to be on the top 8 text chunks based on similarity scores reranked with SentenceTransformerRerank555 https://docs.llamaindex.ai/en/stable/examples/node_postprocessor/SentenceTransformerRerank employing the ‘cross-encoder/ms-marco-MiniLM-L-2-v2’666 https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-2-v2 model in Llamaindex. The evaluator is an external LLM who does not share any document context with the explorer.
@@
(a) Illustration of ‘participant’ annotation under different k values.
(b) Range of ‘participant’ annotation accuracy using LLM-Duo-RAG with GPTs at k=1,2,3.
(c) Token count distribution of retrieved-back chunks across varying similarity scores using ‘participant’ annotation queries of k=1,2,3.
Figure 5: Evaluation of ‘participant’ annotation with POP of different context sizes.
4.2 Case Study: Speech-language Intervention Discovery
Speech-language therapy provides interventions for individuals with speech-language deficits, enhancing their quality of life across various life stages. When choosing an intervention, evidence-based practice (EBP) is attractive as it integrates research evidence from literature into the decision-making process to ensure high-quality patient care Law et al. (1996). Intervention research, especially studies that offer clear intervention frameworks and comprehensive case studies, are valuable references to guide EBP designs. Intervention discovery aims to extensively gather speech-language interventions from the literature corpus as references to facilitate EBP design. It involves identifying relevant studies and extracting essential features of interventions including target disorder, procedure, efficacy, case study, therapy activity, etc., which is extremely labor-intensive for human reviewers, highlighting the efficiency of automating knowledge discovery based on LLMs.
To verify the effectiveness of our method in a realistic scenario, we employ our framework in a speech-language intervention discovery setting. The ontology is shown in Figure 4. To enable a large-scale discovery, we cultivate a literature base including 64,177 papers within the domain of speech-language therapy.
@@ 4.2 Case Study: Speech-language Intervention Discovery
Table 1: Comparison of annotation results with baselines using different LLMs.
4.3 Annotation Baselines
The core idea of our automated intervention framework is to leverage the POP algorithm for guiding the annotation process while utilizing LLM-Duo to refine initial annotations by incorporating external feedback from another LLM. Instead of setting up another LLM for evaluation, recent studies demonstrate that LLMs can engage in self-correction to enhance their responses autonomously Liu et al. (2024)Li et al. (2024b). Notable examples of this include Chain-of-Thought (CoT) Wei et al. (2022) and Self-Refine Madaan et al. (2024). We separately equip the explorer chatbot based on RAG with these two prompting methods for annotation and denote them as CoT-RAG and Self-Refine-RAG. Additionally, in LLM-Duo, a potential substitution of explorer is long-context LLM, which is capable of processing entire document tokens instead of chunking and retrieval with RAG. We refer to the LLM-Duo system as LLM-Duo-RAG when using explorer built on RAG, and as LLM-Duo-LongContext when utilizing long-context LLMs. Besides, we also compare with methods that directly input paper text to LLMs for zero-shot QA annotation without the evaluation feedback loop, including ShortContext LLM, LongContext LLM, OpenAI Assistant, and RAG.
4.4 Evaluation
In the experiment of comparing LLM-Duo with annotation baselines, we report six types of metrics: 1) Consistency Rounds (CR): the number of refine loops the method makes before the annotation generation achieving consistency; 2) Verbosity Index (VI): the number of covered aspects per 1k tokens in the annotations, which is an important metric for emphasizing content completeness; 3) Enumeration Quantity (EQ): the number of items listed in the annotations (i.e., therapy activities, therapy goals.); 4) Faithfulness (Faith): the extent of the annotation faithful to the provided literature literature, which is measured by FaithfulnessEvaluator777 https://docs.llamaindex.ai/en/stable/examples/evaluation/faithfulness_eval of Llamaindex. 5) Accuracy (ACC): the percentage of correct annotations in all LLM-generated annotations. 6) Cover: the percentage of correct LLM-generated annotations to the total mentioned concept entities in the provided literature.
5 Results
In this section, we first provide a detailed evaluation of our progressive ontology prompting algorithm and the LLM-Duo annotation framework. Then, we showcase the results of speech-language intervention discoveries using our automated intervention discovery framework.
5.1 POP Algorithm Study
Context Size Analysis. In the POP algorithm, the context size $k$ determines the diversity and volume of information included in the intervention annotation prompt. To assess the impact of context size on annotation quality, we conducted experiments using various $k$ values to generate intervention annotation prompts for LLM-Duo-RAG. As shown in Figure 5(a), we annotate the ‘participant’ concept for the experiment, which was based on a random selection of 8 speech-language pathology literature.
The annotation accuracy is shown in Figure 5(b). The results indicate that as the context size $k$ increases, annotation accuracy improves significantly, suggesting that a larger context provides more informative prompts, thereby enhancing annotation quality. Moreover, GPT-4-turbo consistently outperforms GPT-3.5-turbo across all $k$ values, demonstrating that more advanced language models can further improve annotation accuracy. Besides, we inspect the text chunks retrieved back by different informative annotation queries based on various $k$ values. We report the range of similarity scores and token count distribution of retrieved-back chunks in Figure 5(c). The similarity score represents the semantically relevancy between retrieved texts to annotation queries. The results show that for $k=1$, the retrieved text chunks generally have low similarity to the query, and the token count decreases as the similarity score increases, leading to lower annotation quality. In contrast, higher $k$ values, especially $k=2$, yield more relevant retrievals. For $k=2$, the token count increases with higher similarity scores, indicating that richer and more relevant content is captured, resulting in improved annotation quality.
Prioritized BFS Analysis. In the POP algorithm, we employ the out-to-in ratio to prioritize neighboring nodes during BFS-based prompt creation and scheduling. This strategy ensures that nodes with more outgoing edges are visited first, allowing them to provide more context for annotating other nodes. For example, one annotation sequence following prioritized POP over the speech-language intervention ontology is ‘TherapyActivity’$\rightarrow$‘TherapyGoal’$\rightarrow$‘Disorder’. In this section, we compare the ‘Disorder’ annotation results using the POP algorithm with and without prioritization, as well as one-shot annotation without using POP, where the entire KG schema is included on a single annotation prompt to extract all triplets. The results are presented in Table 2. We can observe that applying both the POP and prioritized BFS notably enhances annotation performance.
LLM-Duo-RAG GPT3.5-turbo GPT4-turbo
POP 68.18 72.73
POP Prioritized-BFS 77.28 83.20
POP Prioritized-BFS 81.82 86.37
Table 2: Comparison of annotation results with and without the POP and prioritized BFS.
@@ 5.1 POP Algorithm Study
Figure 6: Ablation study and annotated speech-language intervention statistics.
Disorder Interventions Intervention Examples
Aphasia 202 Phonological therapy, Semantic therapy, Syntax Stimulation Program, Melodic Intonation Therapy (MIT), Multimodal Speech Therapy with tDCS, Cross-Language Generalization Therapy (CLGT), Word Learning Paradigm
Autism Spectrum Disorder 178 Personalized Idiom Intervention (PII), Classroom-wide peer tutoring, Idiom Isolation Intervention, Hanen More Than Words, Parent-Mediated Communication-Focused Treatment, Picture Exchange Communication System (PECS)
Dysphagia 129 Swallowing Maneuver Therapy, Focal Vibration Therapy (FVT), Oral Neuromuscular Training and Vibrational Therapy, Prophylactic Swallowing Intervention, High-speed jaw-opening exercise, Palatal Augmentation Prostheses (PAP)
Stuttering 91 Lidcombe Program, Syllable-timed speech, Electronic devices for stuttering, Computer software for stuttering, Bone Conduction Delayed Feedback Therapy, Fluency Techniques and Fear Reduction, Cognitive Behavioral Therapy (CBT)
Phonological Disorder 74 Nonlinear Phonological Intervention Program, Metronome-paced Speech Therapy, Phonological Awareness and Articulatory Training (PAAT), Phonological Meaning Therapy, Motor-based Intervention Approach
AgeGroup Interventions Intervention Examples
Children 745 Pharyngeal flap procedure, National Health Service (NHS) 1-week intensive course, Ultrasound Visual Biofeedback (U-VBF), Intensive Speech Therapy, Community-Based Speech Therapy Models, Early Vocal Intervention, Auditory-Verbal Therapy (AVT)
School-age Children 571 Intensive Speech Therapy, Early Vocal Intervention, APD intervention, Auditory-Verbal Therapy (AVT), Multisensory Stimulation Therapy, Oral Functional Training (OFT), Rhythmic Reading Training (RRT), Rapid Syllable Transition Treatment (ReST)
Older children 366 Semantic Categorization Therapy, Early Vocal Intervention, Rhythmic Reading Training (RRT), Speech Bulb Reduction Program, Intensive speech and language therapy, Peer-Mediated Intervention, Lidcombe Program, Oral Functional Training (OFT)
Preschoolers 347 Early Vocal Intervention, Treatment-as-usual, The Lidcombe Program, Oral Functional Training (OFT), Speech Production Therapy with Reward System, Phonological Interventions and Contrast Therapy, Cycles Phonological Remediation Approach
Adult 319 Pharyngeal flap procedure, Linguistic Retrieval Therapy (LRT), Oral Hydration Intervention, Physiologic Swallowing Therapy, Myofunctional Intervention (OMT), Orthognathic speech therapy, Eye-Tongue Movement Training, Behavioral Voice Therapy
Table 3: Intervention-disorder examples in our discoveries.
5.2 LLM-Duo with Baselines
@@
In this section, we compare LLM-Duo with several advanced baselines using annotations of 8 randomly selected papers from our speech-language literature corpus. The evaluation focuses on three key dimensions: 1) Intervention Recognition (IR), identifying intervention entities within the literature; 2) Intervention Aspect Summary (IAS), annotating the key aspects (e.g., procedure, therapy activity, therapy goals) of the intervention by capturing and summarizing relevant information from the paper; and 3) Intervention Knowledge Completion (IKC), linking interventions to theme classes (e.g., speech awareness, speech articulation, comprehension, foundation skills, etc.) and setting concept nodes (e.g., home, healthcare facilities, schools, teletherapy, etc.). We use human annotators for the IR and IKC tasks to generate gold-standard results for comparison. In the IAS task, we only ask human annotators to tag relevant text fragments related to specific intervention aspects due to potential individual bias in human interpretation.
@@
The experimental results are reported in Table 1. It should be noted that we implemented ‘ShortContext’ using Llama3-instruct-70b (FP16) and Mistral-instruct-8x22b models (INT8). However, directly prompting these models with full paper text fails to produce annotations in a zero-shot QA setting. Their generations do not align with the annotation questions. The results in Table 1 show that LLM-Duo-RAG outperforms all baselines. Although GPT4-turbo has a 128k context window length and is capable of generating annotations, its annotation coverage remains inadequate. Integrating it with the LLM-Duo framework can significantly improve both the accuracy and the comprehensiveness of the intervention annotations. Additionally, compared with simple RAG, self-correct prompting methods such as CoT and Self-Refine can significantly enhance intervention annotations, but their performance is still worse than LLM-Duo-RAG. Instead of utilizing costly GPT models, LLM-Duo-RAG, which employs open-source models including Llama3-instruct-70b and Mistral-instruct-8x22b, can achieve comparable annotation quality.
5.3 Ablation Study
@@
In our implementation, the RAG technique serves as the backbone of explorer. We employ ‘FastCoref’ for coreference resolution and rerank retrieved chunks by similarity score using the ‘cross-encoder/ms-marco-MiniLM-L-2-v2 model’. This section presents ablation studies for both components. We report the accuracy of intervention recognition in this section. As shown in Figure 6(a), the results demonstrate that removing these components significantly decreases annotation accuracy, showing the necessity of each module.
5.4 Speech-Language Intervention Discovery
@@
Through our automated intervention discovery framework, we identified 2,421 interventions supported by case studies from 64,177 literature in the speech-language pathology domain. The statistics of discovered interventions are presented in Figure 6(b) and Figure 6(c). More intervention examples are provided in Table 3. 19 clinicians and students reviewed our annotations through online Google forms. We have constructed the first intervention knowledge graph in the speech-language pathology domain, which will be made publicly accessible upon acceptance. This knowledge graph is anticipated to be a valuable resource for domain experts, facilitating evidence-based clinical decision-making, question-answering, and recommendation systems, ultimately contributing to improved healthcare outcomes.
6 Conclusion
@@
In this paper, we developed a novel LLM-based framework for automatic intervention discovery from literature, featuring a progressive ontology prompting algorithm and a dual-agent system. The proposed method achieves superior performance compared with advanced baselines, enabling more accurate intervention discoveries. Our approach is adaptable to various intervention ontologies in healthcare and offers practical value to improve healthcare quality.
Acknowledgments
The research was supported, in part, by the National AI Institute for Exceptional Education (NSF Award #2229873), Center for Early Literacy and Responsible AI (IES Award #R305C240046), FuSe-TG (NSF Award #2235364) and SaTC (NSF Award #2329704). The opinions expressed are those of the authors and do not represent the views of any sponsors.
Appendix A POP Algorithm
A.1 Algorithm Pseudocode
Algorithm 1 Progressive Ontology Prompting
function OntologyTraversal($\mathcal{G}$, $k$)
    Initialize: $queue,Order\leftarrow[]$, $visited,Khops\leftarrow\{\}$
    Identify the source node $s$
    $enqueue(queue,(s,0))$, $visited[s]\leftarrow\text{true}$, $Khops[s]\leftarrow[]$
    while not $queue$.empty() do
       $(current,hop)\leftarrow dequeue(queue)$
       $Order$.append$(current)$
       Gather unvisited neighbors of $current$ into $neighbors$
       Sort $neighbors$ by out-to-in ratio in descending order
       for $neighbor$ in $neighbors$ do
          if not $visited[neighbor]$ then
             $enqueue(queue,(neighbor,hop+1))$
             $visited[neighbor]\leftarrow\text{true}$
             $Khops[neighbor]\leftarrow[]$
          end if
          if $hop+1\leq k$ then
             $Khops[neighbor].append(current)$
          end if
       end for
    end while
    return $Order$, $Khops$
end function
function Main($\mathcal{G}$, $k$, $T$)
    Initialize $annotations\leftarrow\{\}$
    $Order,Khops\leftarrow\textsc{OntologyTraversal}(\mathcal{G},k)$
    while not $Order$.empty() do
       $current\leftarrow Order$.dequeue(), $context\leftarrow Khops[current]$
       if not $context$.empty() then
          $discoveries\leftarrow\{annotations[t]\text{ for }t\text{ in }context\}$
          $prompt\leftarrow T(current,G_{sub}(context),discoveries)$
       else
          $prompt\leftarrow T(current)$
       end if
       $annotations[current]\leftarrow\textsc{LLM}(prompt)$
    end while
end function
A.2 Annotation Prompt Template Generation
In the POP algorithm, the annotation prompt for a specific concept node is generated from a template $T$, which is constructed using the node’s context (the k-hop visited neighborhood) and the completed annotations within that context. We carefully design the prompt to task LLM to generate $T$ as follows:
Example:
Annotate: Participant
Prompts:
T1: Who is the participant in {#case study}?
Some example output of GPT3.5 for the above prompt is:
T1: What is the frequency used in {#case study}?
T3: {#Intervention} is studied in {#case study} with {#Participant}. What is the frequency of the intervention in {#case study}?
Since template $T$ is composed of concept nodes visited before the current node, the final annotation prompt for the current node is derived by incorporating the annotations associated with these concept nodes into $T$.
Appendix B Speech-Language Intervention Annotation